# A Survey of Personalized Large Language Models: Progress and Future Directions

### v1

@@ Abstract
Large Language Models (LLMs) excel in handling general knowledge tasks, yet they struggle with user-specific personalization, such as understanding individual emotions, writing styles, and preferences. Personalized Large Language Models (PLLMs) tackle these challenges by leveraging individual user data, such as user profiles, historical dialogues, content, and interactions, to deliver responses that are contextually relevant and tailored to each user’s specific needs. This is a highly valuable research topic, as PLLMs can significantly enhance user satisfaction and have broad applications in conversational agents, recommendation systems, emotion recognition, medical assistants, and more. This survey reviews recent advancements in PLLMs from three technical perspectives: prompting for personalized context (input level), finetuning for personalized adapters (model level), and alignment for personalized preferences (objective level). To provide deeper insights, we also discuss current limitations and outline several promising directions for future research. Updated information about this survey can be found at the Github Repo.
@@ 1 Introduction
In recent years, substantial progress has been made in Large Language Models (LLMs) such as GPT, PaLM, LLaMA, DeepSeek, and their variants. These models have demonstrated remarkable versatility, achieving state-of-the-art performance across various natural language processing tasks, including question answering, reasoning, and machine translation (Zhao et al., 2023), with minimal task-specific adaptation.
@@
While LLMs excel in general knowledge and multi-domain reasoning, their lack of personalization creates challenges in situations where user-specific understanding is crucial. For instance, conversational agents need to adapt to a user’s preferred tone and incorporate past interactions to deliver relevant, personalized responses. 8As LLMs evolve, integrating personalization capabilities has become a promising direction for advancing human-AI interaction across diverse domains such as education, healthcare, and finance (Hu et al., 2024; Zhang et al., 2024e, d; Zhu et al., 2024; Wang et al., 2023a, 2024a).
Technical Challenges
Despite its promise, personalizing LLMs presents several challenges. These include efficiently representing and integrating diverse user data, addressing privacy concerns, managing long-term user memories, adaptivity accommodating users’ diverse needs and evolving behaviors (Salemi et al., 2023). Moreover, achieving personalization often requires balancing accuracy and efficiency while addressing biases and maintaining fairness in the generated outputs.
[…]
Personalized Large Language Models
@@
Despite growing interest, the field of PLLMs lacks a systematic review that consolidates recent advancements. This survey aims to bridge the gap by systematically organizing existing research on PLLMs and offering insights into their methodologies and future directions. The contributions of this survey can be summarized as follows: (1) A structured taxonomy: We propose a comprehensive taxonomy, providing a technical perspective on the existing approaches to building PLLMs. (2) A comprehensive review: We systematically review state-of-the-art methods for PLLMs, analyzing fine-grained differences among the methods. (3) Future directions: We highlight current limitations, such as data privacy and bias, and outline promising avenues for future research, including multimodal personalization, edge computing, lifelong updating, trustworthiness, and other emerging challenges.
@@ 2.1 Large Language Models
Large Language Models (LLMs) generally refer to models that utilize the Transformer architecture and are equipped with billions of parameters trained on trillions of text tokens. These models have demonstrated substantial improvements in a myriad of tasks related to natural language understanding and generation, increasingly proving beneficial in assisting human activities. In this work, we mainly focus on autoregressive LLMs, which are based on two main architectures: decoder-only models and encoder-decoder models. Encoder-decoder models such as Flan-T5 (Chung et al., 2022) and ChatGLM (Zeng et al., 2022) analyze input through the encoder for semantic representations, making them effective in language understanding in addition to generation. Decoder-only LLMs focus on left-to-right generation by predicting the next token in a sequence, with numerous instances (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Guo et al., 2025) under this paradigm achieving breakthroughs in advanced capabilities like instruction following and reasoning.
However, these models are typically pre-trained on general-purpose data and lack an understanding of specific user information. As a result, they are unable to generate responses tailored to a user’s unique tastes, preferences, and expectations, limiting their effectiveness in personalized applications where user-specific adaptation is critical.
@@ 2.2 Problem Statement
Personalized Large Language Models (PLLMs) generate responses that align with the user’s style and expectations, offering diverse answers to the same query for different users (Clarke et al., 2024). A PLLM is defined as an LLM that generates responses conditioned not only on an input query , but also on a user ’s personalized data . It aims to predict the most probable response sequence given a query and the personalized context , such that: The personalized data may encapsulate information about the user’s preferences, history, context, and other user-specific attributes. These can include (Figure 1):
@@
Profile/Relationship: User profile, including attributes (e.g., name, gender, occupation), and relationships (e.g., friends, family members), such as .
@@
Historical Dialogues: Historical dialogues, such as question-answer pairs that user interacts with the LLM (e.g., ), where each is a query and is the corresponding answer.
@@
Historical Content: Includes documents, previous reviews, comments or feedback from user . For example, .
@@
Historical Interactions: Includes historical interactions, preferences, ratings from user . For example, .
@@
By incorporating personalized data, PLLMs enhance traditional LLMs, improving response generation, recommendations, and classification tasks.
@@
Note that our survey differs significantly from role-play related LLM personalization (Tseng et al., 2024; Chen et al., 2024a; Zhang et al., 2024f). While role-play focuses on mimicking characters during conversations, PLLMs in this survey focus on understanding users’ contexts and preferences to meet their specific needs. Compared to (Zhang et al., 2024f), which emphasizes broad categories, our work provides a systematic analysis of techniques to enhance PLLM efficiency and performance, with a more detailed technical classification.
@@ 2.3 Proposed Taxonomy
We propose a taxonomy (as illustrated in Figure 1 and Figure 2) from technical perspectives, categorizing the methods for Personalized Large Language Models (PLLMs) into three major levels: (1) Input level: Personalized Prompting focuses on handling user-specific data outside the LLM and injecting it into the model. (2) Model level: Personalized Adaptation emphasizes designing a framework to efficiently fine-tune or adapt model parameters for personalization. (3) Objective Level: Personalized Alignment aims to refine model behavior to align with user preferences effectively. Due to space limitations, analysis papers, datasets, and benchmarks are summarized in the Github Repo.
@@ 3 Personalized Prompting
Prompt engineering acts as a bridge for interaction between users and LLMs. In this survey, prompting involves guiding an LLM to generate desired outputs using various techniques, from traditional text prompts to advanced methods like soft embedding. Soft embedding can be extended not only through input but also via cross-attention or by adjusting output logits, enabling more flexible and context-sensitive responses.
The framework can be expressed as, for each user :
where, is the LLM model that generates the response; is a function that extracts relevant context from the user’s personal context ; represents the combination operator that fuses the query and the relevant personalized context , producing enriched information for the LLM.
@@ 3.1 Profile-Augmented Prompting
Profile-augmented prompting methods explicitly utilize summarized user preferences and profiles in natural language to augment LLMs’ input at the token level ( is the summarizer model). Figure 3(a) shows the illustration.
Non-tuned Summarizer
A frozen LLM can be directly used as the summarizer to summarize user profiles due to its strong language understanding capabilities, i.e., . For instance, Cue-CoT (Wang et al., 2023b) employs chain-of-thought prompting for personalized profile augmentation, using LLMs to extract and summarize user status (e.g., emotion, personality, and psychology) from historical dialogues. PAG (Richardson et al., 2023) leverages instruction-tuned LLMs to pre-summarize user profiles based on historical content. The summaries are stored offline, enabling efficient personalized response generation while meeting runtime constraints. ONCE (Liu et al., 2024b) prompts closed-source LLMs to summarize topics and regions of interest from users’ browsing history, enhancing personalized recommendations.
Tuned Summarizer
Black-box LLMs are sensitive to input noise, like off-topic summaries, and struggle to extract relevant information. Thus, training the summarizer to adapt to user preferences and style is essential. Matryoshka (Li et al., 2024a) uses a white-box LLM to summarize user histories, similar to PAG, but fine-tunes the summarizer instead of the generator LLM. RewriterSlRl (Li et al., 2024b) rewrites the query instead of concatenating summaries, optimized with supervised and reinforcement learning.
CoS (He et al., 2024) is a special case that assumes a brief user profile and amplifies its influence in LLM response generation by comparing output probabilities with and without the profile, adjusting personalization without fine-tuning.
@@ 3.2 Retrieval-Augmented Prompting
Retrieval-augmented prompting (Gao et al., 2023; Fan et al., 2024; Qiu et al., 2024) excels at extracting the most relevant records from user data to enhance PLLMs (See Figure 3(b)). Due to the complexity and volume of user data, many methods use an additional memory for more effective retrieval. Common retrievers including sparse (e.g., BM25 (Robertson et al., 1995)), and dense retrievers (e.g., Faiss (Johnson et al., 2019), Contriever (Izacard et al., 2021)). These methods effectively manage the increasing volume of user data within the LLM’s context limit, improving relevance and personalization by integrating key evidence from the user’s personalized data.
3.2.1 3.2.1 Personalized Memory Construction
@@
Non-Parametric Memory
This category maintains a token-based database, storing and retrieving information in its original tokenized form without using parameterized vector representations. For example, MemPrompt (Madaan et al., 2022) and TeachMe (Dalvi et al., 2022) maintain a dictionary-based feedback memory (key-value pairs of mistakes and user feedback). MemPrompt focuses on prompt-based improvements, whereas TeachMe emphasizes continual learning via dynamic memory that adapts over time. MaLP (Zhang et al., 2024a) further integrates multiple memory types, leveraging working memory for immediate processing, short-term memory (STM) for quick access, and long-term memory (LTM) for storing key knowledge.
Parametric Memory
Recent studies parameterize and project personalized user data into a learnable space, with parametric memory filtering out redundant context to reduce noise. For instance, LD-Agent (Li et al., 2024c) maintains memory with separate short-term and long-term banks, encoding long-term events as parametric vector representations refined by a tunable module and retrieved via an embedding-based mechanism. MemoRAG (Qian et al., 2024), in contrast, adopts a different approach by utilizing a lightweight LLM as memory to learn user-personalized data. Instead of maintaining a vector database for retrieval, it generates a series of tokens as a draft to further guide the retriever, offering a more dynamic and flexible method for retrieval augmentation.
3.2.2 3.2.2 Personalized Memory Retrieval
The key challenge in the personalized retriever design lies in selecting not only relevant but also representative personalized data for downstream tasks. LaMP (Salemi et al., 2023) investigates how retrieved personalized information affects the responses of large language models (LLMs) through two mechanisms: in-prompt augmentation (IPA) and fusion-in-decoder (FiD). PEARL (Mysore et al., 2023) and ROPG (Salemi et al., 2024) similarly aim to enhance the retriever using personalized generation-calibrated metrics, improving both personalization and text quality of retrieved documents. Meanwhile, HYDRA (Zhuang et al., 2024) trains a reranker to prioritize the most relevant information additionally from top-retrieved historical records for enhanced personalization.

### v2

@@ Abstract
Large Language Models (LLMs) excel in handling general knowledge tasks, yet they struggle with user-specific personalization, such as understanding individual emotions, writing styles, and preferences. Personalized Large Language Models (PLLMs) tackle these challenges by leveraging individual user data, such as user profiles, historical dialogues, content, and interactions, to deliver responses that are contextually relevant and tailored to each user’s specific needs. This is a highly valuable research topic, as PLLMs can significantly enhance user satisfaction and have broad applications in conversational agents, recommendation systems, emotion recognition, medical assistants, and more. This survey reviews recent advancements in PLLMs from three technical perspectives: prompting for personalized context (input level), finetuning for personalized adapters (model level), and alignment for personalized preferences (objective level). To provide deeper insights, we also discuss current limitations and outline several promising directions for future research.
Keywords Large Language Models Personalization Memory Text Generation Recommendation
Contents
@@ 1 Introduction
In recent years, substantial progress has been made in Large Language Models (LLMs) such as GPT, PaLM, LLaMA, DeepSeek, and their variants (Zhao et al., 2023). These models have demonstrated remarkable versatility, achieving state-of-the-art performance across various natural language processing (NLP) tasks, including question answering, logical reasoning, and machine translation (Chang et al., 2024; Hu et al., 2024; Zhang et al., 2024a, b; Zhu et al., 2024a; Wang et al., 2023a, 2024a; Li et al., 2024a; Zhao et al., 2025a; Zhu et al., 2025), with minimal task-specific adaptation.
@@
While LLMs excel in general knowledge and multi-domain reasoning, their lack of personalization creates challenges in situations where user-specific understanding is crucial. For instance, conversational agents need to adapt to a user’s preferred tone and incorporate past interactions to deliver relevant, personalized responses. As LLMs evolve, integrating personalization capabilities has become a promising direction for advancing human-AI interaction across diverse domains such as recommendation, education, healthcare, and finance (Wang et al., 2024b; Hu et al., 2024; Zhang et al., 2024a, b; Zhu et al., 2024a; Wang et al., 2023a, 2024a). Despite its promise, personalizing LLMs presents several challenges. These include efficiently representing and integrating diverse user data, addressing privacy concerns, managing long-term user memories, inferring users’ implicit preferences, etc (Salemi et al., 2023). Moreover, achieving personalization often requires balancing accuracy and efficiency while addressing biases and maintaining fairness in the outputs.
@@
Despite growing interest, the field of PLLMs lacks a systematic review that consolidates recent advancements. This survey aims to bridge the gap by systematically organizing existing research on PLLMs and offering insights into their methodologies and future directions. The contributions of this survey are as follows:
A systematic formulation of personalization scenarios: We provide precise problem formulations for different personalized LLM scenarios based on query types and personalized data characteristics, establishing a unified framework that distinguishes between different personalized techniques (Section 2.2).
A structured taxonomy: We propose a comprehensive taxonomy, providing a technical perspective on the existing approaches to building PLLMs across three levels: input-level prompting, model-level adaptation, and objective-level alignment (Section 2.3).
- (3)
A benchmark and evaluation summarization: We provide summarization of metrics, benchmarks tailored to different query types (extraction, abstraction, and generalization) and personalization scenarios (Section 6).
Future directions: We highlight current limitations and outline promising avenues for future research, including multimodal personalization, edge computing, lifelong updating, trustworthiness, etc (Section 7).
@@ 2.1 Large Language Models
Large Language Models (LLMs) generally refer to models that utilize the Transformer architecture and are equipped with billions of parameters trained on trillions of text tokens. These models have demonstrated substantial improvements in a myriad of tasks related to natural language understanding and generation, increasingly proving beneficial in assisting human activities. In this work, we mainly focus on autoregressive LLMs, which are based on two main architectures: decoder-only models and encoder-decoder models. Encoder-decoder models such as Flan-T5 (Chung et al., 2022) and ChatGLM (Zeng et al., 2022) analyze input through the encoder for semantic representations, making them effective in language understanding in addition to generation. Decoder-only LLMs focus on left-to-right generation by predicting the next token in a sequence, with numerous instances (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Guo et al., 2025) under this paradigm achieving breakthroughs in advanced capabilities.
However, these models are typically pre-trained on general-purpose data and lack an understanding of specific user information. As a result, they are unable to generate responses tailored to a user’s unique tastes and expectations, limiting their effectiveness in personalized applications where user-specific adaptation is critical.
@@ 2.2 Problem Statement
Personalized Large Language Models (PLLMs) generate responses that align with the user’s style and expectations, offering diverse answers to the same query for different users (Clarke et al., 2024).
Problem Formulation
Let be a finite set of users. For each :
, where and is the query space;
is the set of ’s personalized data, where denotes the space of all user personalized data.
Let be the output space. For each user and query , there exists a desired output that aligns with ’s preferences and expectations. We define a metric function that evaluates the quality of the LLM-generated output compared with the desired output (the summarization of metrics is shown in Section 6.1, Table 5). Here, we generally assume quantifies the degree of match between the desired and model outputs, with higher values indicating better personalization performance. All notations are summarized in Table 8.
Note that represents the tunable parameters that can exist either within the base LLM architecture itself (when fine-tuning is permitted) or within extra modules incorporated into the LLM framework, like the LoRA module (Yang et al., 2024). Based on personalized data , queries , and outputs (responses) , personalization can be further categorized into multiple scenarios (Figure 2).
Personalized Data
The personalized data may encapsulate information about the user’s preferences, history, context, and other user-specific attributes. These can include profile/relationship, historical dialogues, historical content, and predefined human preferences. A detailed description, including examples, of the classification of input data types for user is provided:
@@
Profile/Relationship User profile, including attributes (e.g., name, gender, occupation), and relationships (e.g., friends, family members), such as .
@@
Historical Dialogues Historical dialogues, such as question-answer pairs that user interacts with the LLM (e.g., ), where each is a query and is the answer.
@@
Historical Content Includes documents, previous reviews, comments or feedback from user . For example, .
@@
Historical Interactions Includes historical interactions, preferences, ratings from user . For example, .
Pre-defined Human Preference: Define a set containing of preference dimensions such as “Helpfulness”. Choose various combinations of these dimensions, form individual preferences, and incorporate them as the instruction. For example, a preference prompt could be “Be harmless and helpful”.
Query Type
Different queries have different focal points regarding the expected responses , which have a certain influence on the method design (high-level comparison with examples is shown in Figure 4 and Table 1). We primarily classify them into the following categories, with further benchmark-based demonstrations provided in Section 6:
Extraction (explicit) This type refers to factual lookup, where the answer can be directly found in the user’s personalized data . The model acts as a retriever:
Here denotes the set of explicit factual tuples contained in . Importantly, extraction also covers simple statistical or aggregative operations (e.g., counting, aggregation) applied over the facts in .
Abstraction (implicit) The model applies a summarization mapping that condenses into a query-independent summary . This abstract profile captures long-term tendencies or preferences, enabling responses beyond explicit tuples:
Generalization (implicit) The model constructs a query-specific state , dynamically combining the user’s data with the current query. The response is then generated from a conditional distribution that also leverages external knowledge :
Unlike , which is global and query-independent, is dynamic and tailored to each query. External knowledge is necessary when alone is insufficient.
| Type | Depends on query? | Depends on external ? | Example |
| Extraction | ✓ | ✗ | “What is my occupation?” Singer; “How many coffees have I bought?” 5 |
| Abstraction | ✗ | ✗ | “Based on the movies I watched, what is my preference?” Sweet & Salty |
| Generalization | ✓ | ✓ | “Rewrite this sentence”; “Recommend a new movie I haven’t seen before” |
Downstream Task
@@
By incorporating personalized data, PLLMs enhance traditional LLMs, improving response generation, recommendation, and classification tasks.
Generation Generation tasks typically involve representing a sequence of strings, such as generating answers for users based on their personalized data and questions or generating content according to the user’s writing style to assist their writing, and so forth (Salemi et al., 2023; Kumar et al., 2024; Zhao et al., 2025d; Au et al., 2025).
@@
Note that our survey differs significantly from role-play related LLM personalization (Tseng et al., 2024; Chen et al., 2024b; Zhang et al., 2024f). While role-play focuses on mimicking characters during conversations (Zhu et al., 2024b; Zhao et al., 2024), PLLMs in this survey focus on understanding users’ contexts and preferences to meet their specific needs. Compared to (Zhang et al., 2024f), which emphasizes broad categories, our work provides a quite different systematic analysis of techniques to enhance PLLM efficiency and performance, with a detailed technical classification.
@@ 2.3 Proposed Taxonomy
We propose a taxonomy (as illustrated in Figure 2 and Figure 3) from technical perspectives (i.e., regarding the personalization operation ), categorizing the methods for Personalized Large Language Models (PLLMs) into three major levels:
Input level: Personalized Prompting focuses on handling user-specific data outside the LLM and injecting it into the model (Section 3).
Model level: Personalized Adaptation emphasizes designing a framework to efficiently fine-tune or adapt model parameters for personalization (Section 4).
Objective Level: Personalized Alignment aims to refine model behavior to align with user preferences effectively (Section 5).
@@ 3 Personalized Prompting
Prompt engineering acts as a bridge for interaction between users and LLMs. In this survey, prompting involves guiding an LLM to generate desired outputs using various techniques, from traditional text prompts to advanced methods like soft embedding. Soft embedding can be extended not only through input but also via cross-attention or by adjusting output logits, enabling more flexible and context-sensitive responses. For each user , the personalization operation is expressed as
where, is a function that extracts relevant context from the user’s personal context ; represents the combination operator that fuses the query and the relevant personalized context , producing enriched information for the LLM. Based on different designs of the operation, we further categorize prompting-based methods into four types, as shown in Figure 5 and Table 2: (1) profile-augmented prompting; (2) retrieval-augmented prompting; (3) soft-fused prompting; and (4) contrastive prompting.
@@ 3.1 Profile-Augmented Prompting
Profile-augmented prompting (Figure 5 (a)) explicitly utilizes summarized user preferences and profiles in natural language to augment LLMs’ input at the token level ( is the summarizer model). The summarizer model is typically distinct from the base LLM (generator).
Non-tuned Summarizer A frozen LLM can be directly used as the summarizer to summarize user profiles due to its strong language understanding capabilities, i.e., . For instance, Cue-CoT (Wang et al., 2023b) employs chain-of-thought prompting for personalized profile augmentation, using LLMs to extract and summarize user status (e.g., emotion, personality, and psychology) from historical dialogues. PAG (Richardson et al., 2023) leverages instruction-tuned LLMs to pre-summarize user profiles based on historical content. The summaries are stored offline, enabling efficient personalized response generation while meeting runtime constraints. ONCE (Liu et al., 2024a) prompts closed-source LLMs to summarize topics and regions of interest from users’ browsing history, enhancing personalized recommendations. DPL (Qiu et al., 2025a) enhances personalization through inter-user comparison, identifying similar users in the same cluster. Leveraging central users from the same cluster where the target user is located creates more accurate preference summaries.
Tuned Summarizer Black-box LLMs are sensitive to input noise, like off-topic summaries, and struggle to extract relevant information. Thus, training the summarizer to adapt to user preferences and style is essential. Matryoshka (Li et al., 2024b) uses a white-box LLM to summarize user histories, similar to PAG, but fine-tunes the summarizer instead of the generator LLM. RewriterSlRl (Li et al., 2024c) rewrites the query instead of concatenating summaries, optimized with supervised and reinforcement learning.
| Profile-Augmented (Section 3.1) | Retrieval-Augmented (Section 3.2) | Soft-Fused (Section 3.3) | Contrastive (Section 3.4) | |
| Summerizer () | Retriever | Encoder | Extractor | |
| Token Combination | Token Combination | Input Prefix; Attention; Logits | Hidden States; Logits | |
| Pros. | Efficiency | Long-Term Memory | Semantic Nuances |
Interpretability
Controllable |
| Cons. | Infromation Loss | Computational Limits Irrelevant Data | Lack of Interpretability | Hyperparameter Sensitivity |
@@ 3.2 Retrieval-Augmented Prompting
Retrieval-augmented prompting (Gao et al., 2023; Fan et al., 2024; Qiu et al., 2024) excels at extracting the most relevant records from user data to enhance PLLMs (See Figure 5 (b)). Due to the complexity and volume of user data, many methods use an additional memory for more effective retrieval (Tan et al., 2025). Common retrievers include sparse (e.g., BM25 (Robertson et al., 1995)), and dense retrievers (e.g., Faiss (Johnson et al., 2019), Contriever (Izacard et al., 2021)). These methods effectively manage the increasing volume of user data within the LLM’s context limit, improving relevance and personalization by integrating key evidence from the user’s personalized data.
3.2.1 Personalized Memory Construction
@@
Non-Parametric Memory This category maintains a token-based database, storing and retrieving information in its original tokenized form without using parameterized vector representations. For example, MemPrompt (Madaan et al., 2022) and TeachMe (Dalvi et al., 2022) maintain a dictionary-based feedback memory (key-value pairs of mistakes and user feedback). MemPrompt focuses on prompt-based improvements, whereas TeachMe emphasizes continual learning via dynamic memory that adapts over time. MaLP (Zhang et al., 2024c) further integrates multiple memory types, leveraging working memory for immediate processing, short-term memory (STM) for quick access, and long-term memory (LTM) for key knowledge. FERMI (Kim and Yang, 2024) maintains misaligned responses and creates personalized prompts for users by using LLMs to progressively refine prompting strategies based on user profiles and past opinions.
Parametric Memory Recent studies parameterize and project personalized user data into a learnable space, with parametric memory filtering out redundant context to reduce noise. For instance, LD-Agent (Li et al., 2024d) maintains memory with separate short-term and long-term banks, encoding long-term events as parametric vector representations refined by a tunable module and retrieved via an embedding-based mechanism. MemoRAG (Qian et al., 2024), in contrast, adopts a different approach by utilizing a lightweight LLM as memory to learn user-personalized data. Instead of maintaining a vector database for retrieval, it generates a series of tokens as a draft to further guide the retriever, offering a more dynamic and flexible method for retrieval augmentation. Inspired by cognitive theories of memory, PRIME (Zhang et al., 2025f) introduces episodic and semantic memory mechanisms to enhance the personalization capabilities of LLMs, and employs a slow-thinking strategy to further improve personalized reasoning.
3.2.2 Personalized Memory Retrieval
The key challenge in the personalized retriever design lies in selecting not only relevant but also representative personalized data for downstream tasks. LaMP (Salemi et al., 2023) investigates how retrieved personalized information affects the responses of large language models (LLMs) through two mechanisms: in-prompt augmentation (IPA) and fusion-in-decoder (FiD). PEARL (Mysore et al., 2023) and ROPG (Salemi et al., 2024) similarly aim to enhance the retriever using personalized generation-calibrated metrics, improving both the personalization and text quality of retrieved documents. Meanwhile, HYDRA (Zhuang et al., 2024) trains a reranker to prioritize the most relevant information additionally from top-retrieved historical records for enhanced personalization. RPM (Kim et al., 2025a) focuses on reasoning ability by extracting user-specific factors from user historical personalized data and creating annotated reasoning paths. At inference, it retrieves these examples to guide reasoning-aligned outputs, strengthening personalized responses. HYDRA and RPM primarily target black-box models where access to the base LLM parameters is restricted.